Call for Papers

Post-JICSLP’96 Workshop and 1996 Compulog Net Area Meeting

on

Parallelism and Implementation Technology for (Constraint) Logic Programming Languages

(in conjunction with JICSLP’96, Bonn, Germany) September 5–6, 1996

This workshop will be focused on both design and implementation of logic and parallel logic programming systems. It is a continuation of the series of workshops on Implementations of Logic Programming Systems, previously held in Budapest (1993), Ithaca (1994), and Portland (1995), and the Compulog Net Area Meetings on Parallelism and Implementation Technologies, previously held in Madrid (1993 and 1994) and Utrecht (1995).

Papers from both academia and industry are invited. Preference will be given to the analysis and description of implemented systems (or currently under implementation) and their associated techniques, problems found in their development or design, and steps taken towards the solution of these problems. Topics include, but are not limited to:

• standard and non–standard sequential implementation schemes (e.g., generalization/modification of WAM, translation to C, etc.);

• implementation of parallel logic programming systems;

• balance between compile–time effort and run–time machinery;

• techniques for the implementation of different declarative programming paradigms based on, or extending, logic programming (e.g., constraint logic programming, concurrent constraint languages, equational–logic languages);

• other implementation–related issues, such as memory management, register allocation, use of global optimizations, etc.

Workshop Format

The workshop will be held in conjunction with the 1996 Joint International Conference and Symposium on Logic Programming, which will take place in Bonn, Germany, from September 2 to September 6, 1996. The workshop will be organized in three sessions:

1. an initial session devoted to presentations of accepted papers, including an invited talk;

2. a session devoted to system demonstrations and description of results obtained on existing implementations; and

3. a final panel session dedicated to discussion of current open problems, drawbacks, and future evolutions.

Paper Submissions

Papers must not exceed 15 pages (approximately 5000 words). The title page should include the name, address, telephone number, and electronic mailing address for each author, as well as a list of keywords. A contact author should also be provided. We invite people interested in giving system demonstrations to submit a short description of the system and the results obtained. We encourage authors to submit papers and system descriptions electronically to

                     compulognet-parimp-meeting@dia.fi.upm.es

The electronic submissions should be in L^{A}TEX or PostScript format. L^{A}TEX style files are available via WWW at

                    http://www.cs.nmsu.edu/lldap/jicslp96.html

or

             http://www.clip.dia.fi.upm.es/Projects/COMPULOG/meeting96

We encourage the use of these formats. If electronic submission is not possible, a copy of the paper should be sent by surface mail to

Manuel Carro / Manuel Hermenegildo Compulog-Net Area Meeting Departamento de inteligencia Artificial Facultad de Informática E-28660 Boadilla del Monte, Madrid Spain

At least one of the authors for each accepted paper is expected to attend the meeting and present the work. The collection of accepted papers will be made available at the workshop and published as a UPM/Compulog-Net Technical Report. The papers will also be available electronically after the workshop at the aforementioned WWW addresses.

A limited number of grants to support attendance at the meeting are available. Priority will be given to authors of accepted papers. If you are interested, please send your request to

                     compulognet-parimp-meeting@dia.fi.upm.es

This meeting is sponsored by the Compulog Network on Computational Logic. Attendance is however in no way limited to network members.

Important dates

Deadline for submission:  June 7, 1996
Notification of acceptance: July 1, 1996
Camera-ready papers due: July 26, 1996
Workshop:  September 5 or 6, 1996

Important note: The exact dates and length of the meeting depend largely on the number of papers received and the number of people expected to attend it. Thus, we encourage anyone interested in attending the workshop to submit a note by electronic mail to the workshop organizers

before July, 1, 1996, at the address compulognet-parimp-meeting@dia.fi.upm.es.

Organizing Committee

Manuel Carro, Manuel Hermenegildo  Enrico Pontelli
Departamento de Inteligencia Artificial  Laboratory for Logic and Databases
Facultad de Informática  New Mexico State University
Universidad Politécnica de Madrid  Box 30001, Dept. CS
E–28660 Boadilla del Monte, Madrid, Spain  Las Cruces, NM 88003, U.S.A.
Voice: +34-1-336-7448  Voice: (505)646-6239
Fax: +34-1-352-4819  Fax: (505)646-1002

{mcarro, herme}@fi.upm.es                  epontell@cs.nmsu.edu

Information and Requests

E-mail address:  compulognet-parimp-meeting@dia.fi.upm.es
WWW addresses: http://www.clip.dia.fi.upm.es/Projects/COMPULOG/meeting96

                  http://www.cs.nmsu.edu/lldap/jicslp96.html

Surface mail:  Manuel Carro / Manuel Hermenegildo

Compulog-Net Area Meeting Facultad de Informática — UPM E-28660 Boadilla del Monte, Madrid Spain

Fax:  +34-1-352-4819

JICSLP’96:  http://www.informatik.uni-bonn.de/~jicslp96